A manufacturer of polygraphs lie detector claims that its machines can corrrectly distinguish truthful responses to question 78% of the time if the machine is tested using a set of 60 questions determine the mean and standard deviation for this bernoulli process



1
Expert's answer
2021-11-21T17:13:57-0500

Bernoulli proccess is a sequence of independent equally distributed bernoulli random variables

Let X be a random variable represent the amount of correctly distinguished truthful responses, then

"X=X{\scriptscriptstyle 1}+...+X{\scriptscriptstyle 60}" , where each "X{\scriptscriptstyle 1},...,X{\scriptscriptstyle 60}" can take value 1 with p = 0.78 and 0 with q = 0.22

Which means that X has binominal distribytion with n = 60, p = 0.78

X ~ Bin(0.78, 60)

The mean of X is "m=0.78 * 60 = 46.8"

The variance is "d = 60 * 0.78 * (1-0.78) = 10.296"

The standard deviation is "\sigma = \sqrt{d}=\sqrt{10.296}\approx3.21"